How do I begin to solve this equation (absolute value bar)2x+1/3x+4(absolute value bar) = 1
Answer by stanbon(57361)

How do I begin to solve this equation (absolute value bar)2x+1/3x+4(absolute value bar) = 1
------------------------
|(2x+1)/(3x+4)| = 1
----
Draw a number line.
Plot x = -1/2 and x = -4//3 on the line
----
Determine the sign of the fraction in each of the three line intervals.
------
If x < -4/3 the fraction is -/- = positive
So (2x+1)/(3x+4) = 1
2x+1 = 3x+4
x = -3
----
If -4/3 < x < -1/2, the fraction is -/+ = negative
So (2x+1)/(3x+4) = -1
2x+1 = -3x-4
5x = -5
x = -1
-------------
If x > -1/2 the fraction is +/+ = positive
You get x = -3 as you did above.
====================================
Cheers,
Stan H.